OpenStudy (anonymous):

please help: A particle moves along a straight line and its position at time t is given by s(t)= 2t^3 - 27 t^2 + 84 t where s is measured in feet and t in seconds. what is the TOTAL distance the particle travels between time 0 and time 18

OpenStudy (turingtest):

To solve this problem we need to take notice of when the particle changes direction. This happens when s'(t)=v=0, so s'(t)=v=6t^2-54t+84=0=t^2-9t+14 (t-7)(t-2)=0 so the particle changes direction at t={2,7} Now we must integrate within these intervals separately and take absolute values to avoid negative answers. The intervals are \[0 \le t \le 2\]\[2 \le t \le 7\] and \[7 \le t \le 18\] So integrate\[\int\limits\limits v(t)dt=\int\limits\limits s'(t)dt\] for each interval and take the absolute value of any negative answers. This will give you your solution.

OpenStudy (turingtest):

PS: You must add the absolute value of the results of each of the three integrals to find the TOTAL distance, in case that wasn't obvious.

OpenStudy (anonymous):

thank you so much for your time, would there by another way to solve without using integration?

OpenStudy (turingtest):

hmmm.... I guess if you just integrate\[\int\limits_{a}^{b} s'(t)dt=s(t)\]evaluated from a to b. So it looks like you could just find the intervals and do \[\left| s(2)-s(0) \right|+\left| s(2)-s(7) \right|+\left| s(18)-s(7) \right|=\]and that answer should be the same. Never thought about it that way though.

OpenStudy (turingtest):

sorry, the expression should be |s(2)−s(0)|+|s(7)−s(2)|+|s(18)−s(7)|= but because it's absolute value it doesn't even matter!!! Ia-bI=Ib-aI

OpenStudy (anonymous):

thank you so much, you helped me a bunch and made it really clear =)

OpenStudy (turingtest):

Anytime